OpenStudy (anonymous):

Does anyone know why Pythagorean theorem can be applied on calculating velocity,after all velocity is not length

OpenStudy (anonymous):

I mean why \[V ^{2}=vx ^{2}+vy ^{2}\] Since it is not length we cannot apply Pythagorean theorem

OpenStudy (anonymous):

velocity is a vectorel quantity so whem you add their x and y components you must use vectorel adding.above formula gives you only magnitude not direction but velocity means both magnitude and direction

OpenStudy (irishboy123):

Pythagoreas [not his invention, it has been known for a lot longer] is the most amazing way (think about it) of adding, or aggregating, things that move in different directions or dimensions, viz at right angles. it can be distance (displacement), or speed (velocity), or acceleration, or even space-time (just add in (ct)^2). it also works in the 3rd dimensions. s^2 = x^2 + y^2 + z^2. OR more accurately: s^2 = x^2 + y^2 + z^2 + (ct)^2.

OpenStudy (whpalmer4):

When you apply Pythagoras to \(V^2 = vx^2 + vy^2\) you can consider them instead to be distances covered in unit time. There's a 1:1 mapping between the quantity being calculated and the one to which you apply the PT.